### v2

@@ 4.2 Partial Identification of Causal Effects
In this section, we lay out conditions and derive bounds on average treatment effects. We assume that the researcher has already identified the set of local preferences at a cutoff of interest. This means that the researcher has a set-valued variable $\bm{\mathrm{Q}}_{j}$ for all students in the vicinity of a cutoff $j$ corresponding to a comparable pair $(j,k)$ in $\mathcal{P}$. Researchers may construct $\bm{\mathrm{Q}}_{j}$ using the methods in Section 4.1 if they find it reasonable to rely on at least some of the specific assumptions in that subsection; otherwise, they may use any other method to construct $\bm{\mathrm{Q}}_{j}$. There is no restriction on the choice of the method for constructing $\bm{\mathrm{Q}}_{j}$ except a couple of high-level conditions that we assume to hold in this section. We start by defining the conditional support of partially identified true local preferences.
@@
Assumption 6(i) says that $\bm{\mathrm{Q}}_{j}(\omega)$ of individual $\omega$ contains the true pair of local preferences $Q_{j}(\omega)$ of that individual (for almost all individuals), which is a minimum requirement for the construction of $\bm{\mathrm{Q}}_{j}(\omega)$. This does not say anything about the sharpness of $\bm{\mathrm{Q}}_{j}$. For example, $\bm{\mathrm{Q}}_{j}=\mathcal{J}^{0}\times\mathcal{J}^{0}$ is completely uninformative and trivially satisfies Assumption 6(i). The sharpness requirement is stated in Assumption 6(ii). It says that all possibilities of local preferences listed in $\bm{\mathrm{Q}}_{j}$ actually occur in the data with positive probability. This rules out unnecessarily large sets $\bm{\mathrm{Q}}_{j}$. Assumption 6(ii) may be dropped at the cost of our lacking sharply identified sets in the rest of this section.
@@
    $\mathbb{P}\left[Q_{j}\in A|S_{j}=c_{j}\right]\geq\mathbb{P}\left[\bm{\mathrm{Q}}_{j}\subseteq A|S_{j}=c_{j}^{+}\right];\text{ or }$
@@
Example (SD Example, Continued).
Continue to assume that the four schools are acceptable for everyone. Suppose for a moment that all combinations of $(P,Q)$ that satisfy Assumptions 35 exist in the economy, for both $S_{1}\geq c_{4}$ and $S_{1}<c_{4}$. Then, this is the list of all possible $\bm{\mathrm{Q}}_{4}$:
@@
Recall that the construction of the random set $\bm{\mathrm{Q}}_{j}$ depends on assumptions regarding the behavior of agents when they submit $P$. For example, Section 4.1 characterizes $\bm{\mathrm{Q}}_{j}$ by assuming weak partial order and cutoff characterization wrt $P$. Alternatively, the identification approach of agarwal2018 makes different types of assumptions on agents’ expectations and requires data variation in the choice environment. The theoretical credibility of these types of assumptions depends on the mechanism faced by agents; in practice, the assumptions have testable implications for what we should observe in the data. It is therefore useful to characterize a falsification test based on these implications to aid researchers in screening out assumptions rejected by the data. Such a test can rely on the fact that the right-hand sides of the inequalities in Proposition 4 must provide a lower bound for a probability mass function if the model assumptions are correct. For any partition $\mathcal{A}$ of $\mathcal{J}^{0}\times\mathcal{J}^{0}$, we must have $\sum_{A\in\mathcal{A}}\mathbb{P}\left[Q_{j}\in A|S_{j}=c_{j}\right]=1$ for any given distribution in the sharp set of Proposition 4. Thus, the same sum applied to the right-hand sides of the inequalities above must be less than or equal to one.
Corollary 2 (Model Falsification Test).
@@
[…]
@@
Example (SD Example, Continued).
@@
The following result utilizes the proportions $\underline{\delta}_{j,k}^{+}$ and $\underline{\delta}_{j,k}^{-}$ to partially identify average outcomes for individuals with $Q_{j}=(j,k)$ on either side of the cutoff. Taking differences of these bounds yield bounds for the averages of the treatment effects $Y(j)-Y(k)$.
@@
Suppose Assumptions 2, 6, and 7 hold. Consider a pair $(j,k)\in\mathcal{P}$ such that $\underline{p}_{j,k}>0$.
  1. (i)
    If $g(Y)$ is a continuous random variable for $g\in\mathcal{G}$ of Assumption 2, then we have the following bounds on $\mathbb{E}[g(Y(j))|Q_{j}=(j,k),S_{j}=c_{j}]$ and $\mathbb{E}[g(Y(k))|Q_{j}=(j,k),S_{j}=c_{j}]$:
    $\mathbb{E}\left[g(Y)\left|\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset,g(Y)<F_{j,k+}^{-1}(\underline{\delta}_{j,k}^{+}),S_{j}=c_{j}^{+}\right.\right]$
    $\hskip 28.45274pt\leq~{}\mathbb{E}\left[g(Y(j))\left|Q_{j}=(j,k),S_{j}=c_{j}\right.\right]~{}\leq$
    $\mathbb{E}\left[g(Y)\left|\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset,g(Y)>F_{j,k+}^{-1}(1-\underline{\delta}_{j,k}^{+}),S_{j}=c_{j}^{+}\right.\right],$
@@
    $\mathbb{E}\left[g(Y)\left|\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset,g(Y)<F_{j,k-}^{-1}(\underline{\delta}_{j,k}^{-}),S_{j}=c_{j}^{-}\right.\right]$
    $\hskip 28.45274pt\leq~{}\mathbb{E}\left[g(Y(k))\left|Q_{j}=(j,k),S_{j}=c_{j}\right.\right]~{}\leq$
    $\mathbb{E}\left[g(Y)\left|\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset,g(Y)>F_{j,k-}^{-1}(1-\underline{\delta}_{j,k}^{-}),S_{j}=c_{j}^{-}\right.\right],$
    where $F_{j,k+}^{-1}(u):=\inf\left\{y:\mathbb{P}\left[g(Y)\leq y\left|\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset,S_{j}=c_{j}^{+}\right.\right]\geq u\right\}$ and
@@
  2. (ii)
    If $Y$ is a binary random variable, then we have the following bounds on $\mathbb{E}[Y(j)|Q_{j}=(j,k),S_{j}=c_{j}]$ and $\mathbb{E}[Y(k)|Q_{j}=(j,k),S_{j}=c_{j}]$:
    $\hskip 28.45274pt\leq~{}\mathbb{E}\left[Y(j)\left|Q_{j}=(j,k),S_{j}=c_{j}\right.\right]~{}\leq$
    and
    $\hskip 28.45274pt\leq~{}\mathbb{E}\left[Y(k)\left|Q_{j}=(j,k),S_{j}=c_{j}\right.\right]~{}\leq$
The bounds in Proposition 5 build on the work by horowitz1995. To see the intuition, take part (i) of the proposition, make $g(Y)=Y$, and focus on individuals just above the cutoff. Among all individuals in the subpopulation with $\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset$ and $S_{j}=c_{j}^{+}$, there are ${\delta}^{+}_{j,k}$ of them who have $Q_{j}=(j,k)$ and $Y=Y(j)$ by the cutoff characterization. We do not know who these individuals are among those in the subpopulation. However, the lowest possible value for $\mathbb{E}[Y(j)|Q_{j}=(j,k),S_{j}=c_{j}]$ occurs if all such individuals are located at the lower tail of the distribution of outcomes in the subpopulation. Likewise, the highest possible value for $\mathbb{E}[Y(j)|Q_{j}=(j,k),S_{j}=c_{j}]$ occurs if all of that same fraction of individuals are located in the upper tail of the distribution of outcomes. We do not know ${\delta}^{+}_{j,k}$, but we do know that it is no smaller than $\underline{\delta}^{+}_{j,k}>0$. The bounds only grow wider as the fraction ${\delta}^{+}_{j,k}$ decreases, so the bounds evaluated at ${\delta}^{+}_{j,k}=\underline{\delta}^{+}_{j,k}$ take into account all possible values for ${\delta}^{+}_{j,k}$. The expressions for the bounds in the case that $g(Y)=Y$ is binary change relative to $g(Y)$ continuously, but the derivation of the bounds follows the same intuition. We refer the reader to the proof in the appendix for details (Section A.8).
Although intuitive, these bounds are not necessarily sharp because $\underline{\delta}^{+}_{j,k}$ is not exogenously given as in horowitz1995; $\underline{\delta}^{+}_{j,k}$ is constructed from only the distribution of $Q_{j}$. Providing a complete characterization of the sharp bounds is complex because it involves deriving bounds on the joint distribution of potential outcomes and $Q_{j}$, which may not be practical when the potential outcomes are continuous. For the sake of simplicity, we relegate the sharp characterization to Section B in the appendix. The bounds of Proposition 4 contain the sharp bounds of Section B as long as our model assumptions are true. The lack of sharpness may not matter in practice when Proposition 4 yields tight bounds for a given dataset. However, if our assumptions are not true, the bounds of Proposition 4 may not contain the sharp bounds of Section B. That said, we may obtain tight bounds from the data using Proposition 4, but this does not mean they contain the true parameter (kedagni2020). Therefore, it is advisable to assess the testable implications of Corollary 2 as a matter of routine.
5 Assignment to College Majors and Graduation in Chile
@@
5.1 Data, Institutional Setting, and Evidence of Strategic Behavior
@@
We use publicly available data on Chile’s centralized college application and assignment system from 2004 to 2010 and on graduations from 2007 to 2020. The institutional setting has been described in detail by hastingsneilsonzimmerman2013 and larroucaurios2020; larroucaurios2021, among others. College choice in Chile is organized as a semicentralized system—a subset of universities participate in a centralized market in which a clearinghouse collects ranked-order lists from applicants and determines assignments using a variant of the DA algorithm. Students can submit ranked-order lists of up to eight major–university pairs (“programs”) out of more than 1,000.999The cap was increased to ten in 2012. Priorities are program-specific and determined by a weighted average of scores obtained in a national standardized test (the PSU, for prueba de selección universitaria) and of high-school GPA. [^10] In 2014, students’ relative rank within their high school was added as one of the “primary” scores to be averaged to construct priorities. Descriptive statistics about the sample of students and programs are shown in the left panel of Tables 3 and 4 in Appendix C.
@@
That Chilean college applicants behave strategically has been thoroughly documented by larroucaurios2020; larroucaurios2021. Using a 2014 survey linked to administrative data on applications, they show that listed programs often do not coincide with the truly preferred programs explicitly elicited by the survey. Focusing on applications to medicine programs, they find, for instance, that “among the 40,000 students who answered the survey, close to 10% (3,797) reported Medicine as their top preference, and 2,987 of these students ended up applying to the system. Among these, only 1,360 listed Medicine as their top preference.” Moreover, students’ probability of not including medicine at the top of their submitted list (while declaring it their most preferred program in the survey) increases as their application scores decrease, and the probability of applying to medicine drops after the 600–750 application score range, where the cutoffs for most medicine programs lie. This suggests that students tend to omit medicine as their admission chances fall, despite preferring it over other programs. These findings are confirmed in a 2019 survey, where the researchers find that “most students who did not include their first true preference in their application list expected a higher cutoff than the cutoff for their first listed preference.”
@@
Figure 1 provides additional evidence of strategic behavior. Consider a program $j$ with cutoff $c_{j}$. Suppose that students tend to prefer programs of higher quality, consistent with what is found in the literature. If applicants behave strategically, one would expect applications to program $j$ to peak among students with application scores close to $c_{j}$. If cutoffs tend to remain in the same neighborhood across years, students with application scores much higher than $c_{j}$ can expect to be admissible to more selective, higher-quality programs than $j$, which they prefer over $j$. Hence, we expect very few of these students to include $j$ on their list. As application scores drop and are closer to $c_{j}$, students’ chances of admission to the most selective programs decrease, and program $j$ becomes one of the most selective (desirable) programs among those for which they still have a high admission probability. Hence, we expect applications to $c_{j}$ to increase as application scores decrease and draw closer to $c_{j}$. As application scores decrease below $c_{j}$, students realize that their probability of admission to program $j$ is lower, and while program $j$ remains a relatively desirable (selective) alternative, we expect these expectations to drive applications down. This application pattern, expected if students behave strategically, is exactly what we observe in the top panel of Figure 1. Pooling all programs $j$ together, the top panel of Figure 1 shows the fraction of students listing program $j$ in their ranked-order lists (ROLs or $P$ in terms of our notation), as a function of the distance between their priority score for program $j$ and the cutoff $c_{j}$.
@@
It may be difficult to disentangle the role of preferences from the role of expectations about admission probabilities when both may enter students’ choice of which programs to include in their ROLs (manski_ecta2004; agarwal2018). The pattern observed in the top panel of Figure 1 could, alternatively, be consistent with students not behaving strategically but preferring programs that are a good fit in terms of quality, that is, programs in which their skill level would be close to the average skill level. If this were the case, applications should peak among students whose skills (proxied by application score) are close to the mean skill level in the program. This is not what we observe in the bottom panel of Figure 1. Pooling all programs $j$ together, the bottom panel of Figure 1 shows the fraction of students including program $j$ in their list, as a function of the distance between their priority score for program $j$ and the mean application score among students admitted to $j$. Conditional on application score, the share of students applying to a program $j$ is not the highest for application scores close to the mean score among students admitted to $j$. It peaks well below this level, showing that students do not systematically prefer programs in which they would be the “average” student. This is again consistent with the hypothesis that students behave strategically.
@@
This figure provides evidence of strategic behavior in student applications. Pooling all programs $j$ together, the top panel of the figure shows the fraction of students listing program $j$ in their ranked-order lists (ROLs or $P$ in terms of our notation), as a function of the distance between their priority score for program $j$ and the cutoff $c_{j}$. Pooling all programs $j$ together, the bottom panel shows the fraction of students including program $j$ in their list, as a function of the distance between their priority score for program $j$ and the mean application score among students admitted to $j$.

### v3

@@ 4.2 Partial Identification of Causal Effects
In this section, we lay out conditions and derive bounds on average treatment effects. We assume that the researcher has already identified the set of local preferences at a cutoff of interest. This means that the researcher has a set-valued variable $\bm{\mathrm{Q}}_{j}$ for all students in the vicinity of a cutoff $j$ corresponding to a comparable pair $(j,k)$ in $\mathcal{P}$. Researchers may construct $\bm{\mathrm{Q}}_{j}$ using the methods in Section 4.1 if they find it reasonable to rely on at least some of the specific assumptions in that subsection; otherwise, they may use any other method to construct $\bm{\mathrm{Q}}_{j}$. There is no restriction on the choice of the method for constructing $\bm{\mathrm{Q}}_{j}$ except for a couple of high-level conditions that we assume to hold in this section. We start by defining the conditional support of partially identified true local preferences.
@@
Assumption 6(i) says that $\bm{\mathrm{Q}}_{j}(\omega)$ of individual $\omega$ contains the true pair of local preferences $Q_{j}(\omega)$ of that individual (for almost all individuals), which is a minimum requirement for the construction of $\bm{\mathrm{Q}}_{j}(\omega)$. This does not say anything about the sharpness of $\bm{\mathrm{Q}}_{j}$. For example, $\bm{\mathrm{Q}}_{j}=\mathcal{J}^{0}\times\mathcal{J}^{0}$ is completely uninformative and trivially satisfies Assumption 6(i). The sharpness requirement is stated in Assumption 6(ii). It says that all possibilities of local preferences listed in $\bm{\mathrm{Q}}_{j}$ actually occur in the data with positive probability. This rules out unnecessarily large sets $\bm{\mathrm{Q}}_{j}$. Assumption 6(ii) may be dropped at the cost of lacking sharpness in the identified sets in the rest of this section.
@@
[…]
@@
Example (SD Example, Part V).
Continue to assume that the four schools are acceptable for everyone. Suppose for a moment that all combinations of $(P,Q)$ that satisfy Assumptions 35 exist in the economy, for both $S_{1}\geq c_{4}$ and $S_{1}<c_{4}$. Then, the list of all possible $\bm{\mathrm{Q}}_{4}$ is as follows:
@@
Recall that the construction of the random set $\bm{\mathrm{Q}}_{j}$ depends on assumptions regarding the behavior of agents when they submit $P$. For example, Section 4.1 characterizes $\bm{\mathrm{Q}}_{j}$ by assuming weak partial order and cutoff characterization wrt $P$. Alternatively, the identification approach of Agarwal and Somaini (2018) makes different types of assumptions on agents’ expectations and requires data variation in the choice environment. The theoretical credibility of these types of assumptions depends on the mechanism faced by agents; in practice, the assumptions have testable implications for what we should observe in the data. It is therefore useful to characterize a falsification test based on these implications to aid researchers in screening out assumptions rejected by the data. Such a test can rely on the fact that the right-hand sides of the inequalities in Proposition 4 must provide a lower bound for a probability mass function if the model assumptions are correct. For any partition $\mathcal{A}$ of $\mathcal{J}^{0}\times\mathcal{J}^{0}$, we must have $\sum_{A\in\mathcal{A}}\mathbb{P}\left[Q_{j}\in A|S_{j}=c_{j}\right]=1$ for any given distribution in the sharp set of Proposition 4. Thus, the same sum applied to the right-hand sides of the inequalities above must be less than or equal to one.
Corollary 2 (Model’s Falsification Test).
@@
$\sum_{A\in\mathcal{A}}\Bigg{\{}\;\;\;$ $\mathbb{I}\left\{A\in\bm{\mathrm{\Lambda}}^{\cup+}_{j}\cap\bm{\mathrm{\Lambda}}^{\cup-}_{j}\right\}\max\left\{~\mathbb{P}\left[\bm{\mathrm{Q}}_{j}\subseteq A|S_{j}=c_{j}^{+}\right]~;~\mathbb{P}\left[\bm{\mathrm{Q}}_{j}\subseteq A|S_{j}=c_{j}^{-}\right]~\right\}$
@@
Example (SD Example, Part VI).
@@
The following result utilizes the proportions $\underline{\delta}_{j,k}^{+}$ and $\underline{\delta}_{j,k}^{-}$ to partially identify average outcomes for individuals with $Q_{j}=(j,k)$ on either side of the cutoff. Taking differences of these bounds yield bounds for the averages of the treatment effects Y(j)Y(k)Y(j)-Y(k).
@@
Suppose Assumptions 2, 6, and 7 hold. Consider a pair $(j,k)\in\mathcal{P}$ such that $\underline{p}_{j,k}>0$, then we have the following bounds on $\mathbb{E}[g(Y(j))|Q_{j}=(j,k),S_{j}=c_{j}]$ and $\mathbb{E}[g(Y(k))|Q_{j}=(j,k),S_{j}=c_{j}]$:
[…]
@@
[…]
where $U\sim\text{Uniform}[0,1]$ (independent of everything else),
$F_{j,k+}^{-1}(u):=\inf\left\{y:\mathbb{P}\left[g(Y)\leq y\left|\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset,S_{j}=c_{j}^{+}\right.\right]\geq u\right\}$ and
@@
The expressions for the bounds simplify significantly in the cases where $Y$ is binary or continuous. For the sake of brevity, we relegate the detailed formulas to the Appendix (Section B.7). The bounds in Proposition 5 build on the work by Horowitz and Manski (1995). To see the intuition, consider the case where $g(Y)=Y$ is a continuous random variable, and focus on individuals just above the cutoff. Among all individuals in the subpopulation with $\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset$ and $S_{j}=c_{j}^{+}$, a fraction ${\delta}^{+}_{j,k}$ of them has $Q_{j}=(j,k)$ and $Y=Y(j)$ by the cutoff characterization. We do not know who these individuals are among those in the subpopulation. However, the lowest possible value for $\mathbb{E}[Y(j)|Q_{j}=(j,k),S_{j}=c_{j}]$ occurs if all such individuals are located at the lower tail of the distribution of outcomes in the subpopulation. Likewise, the highest possible value for $\mathbb{E}[Y(j)|Q_{j}=(j,k),S_{j}=c_{j}]$ occurs if all of that same fraction of individuals are located in the upper tail of the distribution of outcomes. We do not know ${\delta}^{+}_{j,k}$, but we do know that it is no smaller than $\underline{\delta}^{+}_{j,k}>0$. The bounds only grow wider as the fraction ${\delta}^{+}_{j,k}$ decreases, so the bounds evaluated at ${\delta}^{+}_{j,k}=\underline{\delta}^{+}_{j,k}$ take into account all possible values for ${\delta}^{+}_{j,k}$.
Although intuitive and analytically simple, these bounds are not necessarily sharp because $\underline{\delta}^{+}_{j,k}$ is not exogenously given as considered by Horowitz and Manski (1995); $\underline{\delta}^{+}_{j,k}$ is constructed from the marginal distribution of $Q_{j}$ but there could be additional identifying information in the joint distribution of $Q_{j}$ and $Y$. Providing a complete characterization of the sharp bounds is complex because it involves deriving bounds on the joint distribution of potential outcomes and $Q_{j}$, which may not be practical when the potential outcomes are continuous. For the sake of simplicity, we relegate the sharp characterization to Section A.3 in the appendix. The bounds of Proposition 5 contain the sharp bounds of Section A.3 as long as our model assumptions are true. The lack of sharpness may not matter in practice when Proposition 5 yields tight bounds for a given dataset. However, if our assumptions are not true, the bounds of Proposition 5 may not contain the sharp bounds of Section A.3. That said, we may obtain tight bounds from the data using Proposition 5, but this does not mean they contain the true parameter (Kédagni et al., 2020). Therefore, it is advisable to assess the testable implications of Corollary 2 as a matter of routine.
5 Assignment to College Programs and Graduation in Chile
@@
5.1 Institutional Setting and Evidence of Strategic Behavior
@@
We use publicly available data on Chile’s centralized college application and assignment system from 2004 to 2010 and on graduation from 2007 to 2020. The institutional setting has been described in detail by Hastings et al. (2013) and Larroucau and Rios (2020, 2021), among others. College choice in Chile is organized as a semicentralized system—a subset of universities participate in a centralized market in which a clearinghouse collects rank-order lists from applicants and determines assignments using a variant of the DA algorithm. Students can submit rank-order lists of up to eight major–university pairs (“programs”) out of more than 1,000.101010The cap was increased to ten in 2012. Priorities are program-specific and determined by a weighted average of scores obtained in a national standardized test (the PSU, for prueba de selección universitaria) and of high-school GPA. [^11] In 2014, students’ relative rank within their high school was added as one of the “primary” scores to be averaged to construct priorities. Descriptive statistics about the sample of students and programs are shown in the left panel of Tables C-1 and C-2 in Appendix C.
@@
Larroucau and Rios (2020, 2021) thoroughly document that Chilean college applicants behave strategically. Using a 2014 survey linked to administrative data on applications, they show that listed programs often do not coincide with the truly preferred programs as elicited by the survey. Focusing on applications to medicine programs, they find, for instance, that “among the 40,000 students who answered the survey, close to 10% (3,797) reported Medicine as their top preference, and 2,987 of these students ended up applying to the system. Among these, only 1,360 listed Medicine as their top preference.” Moreover, students’ probability of not including medicine at the top of their submitted list (while declaring it their most preferred program in the survey) increases as their application scores decrease, and the probability of applying to medicine drops after the 700–750 application score range, where the cutoffs for most medicine programs lie. This suggests that students tend to omit medicine as their admission chances fall, despite preferring it over other programs. These findings are confirmed in a 2019 survey, where the researchers find that “most students who did not include their first true preference in their application list expected a higher cutoff than the cutoff for their first listed preference.”
Expectations about cutoffs are a key factor behind an agent’s decision of which programs to list. Students in Chile can easily form expectations by learning from historical data on program cutoffs. In fact, many programs have relatively stable cutoffs over time (see, e.g., Kapor et al. (2024), Figure 1). Individuals know their priority scores before they decide on their submission $P$. Thus, stable cutoffs translate into a relatively high certainty about the set of schools that are feasible to the individual.
Figure 1 shows that more selective programs have not only higher cutoffs but also more stable cutoffs over time. Panel (a) plots the histogram of cutoff values pooled across years and then compares that with two subsets of programs: first, with programs that we select to illustrate our methods in Section 5.2 below; and second, with medicine programs. Section 5.2 focuses on programs with a sufficiently large number of applications because we need adequate sample sizes to implement our estimation procedures. We refer to them as programs of interest. Programs of interest are more selective than most, and Medicine programs stand out as exceptionally selective. We see in Panel (a) that cutoffs generally increase with selectivity and tend to be more compressed at the top of the distribution. In order to view time variability as a share of cutoff value within each program, we divide the standard deviation of cutoffs across time by the average of cutoffs across time. Panel (b) displays the distribution of these program-specific ratios across all programs. That distribution is again compared with the programs of interest of Section 5.2 and with Medicine programs. We find that selectivity is generally associated with stability of cutoffs. For most programs of interest, the standard deviation of cutoff variability is at most 3% of cutoff value, while for medicine programs is at most 1.65%. The survey data evidence on strategic behavior regarding Medicine is consistent with agents having strong a priori knowledge about ex post cutoffs for Medicine.
Figure 1: Cutoffs of Selective Programs
This figure displays distributions of cutoffs of college-program pairs in Chile for the years of 2004–2010. Histograms in both panels show distributions for three sets of programs: all programs (light gray), the subset of programs selected for the analysis of Section 5.2, i.e., programs of interest (medium gray), and medicine programs (dark gray). Panel (a) plots the cutoff values pooled over all years. Panel (b) displays the distribution of program-specific ratios: the standard deviation across the years of a program cutoff divided by the average across the years of the same program cutoff.
@@
Figure 2 provides additional evidence of strategic behavior arising from a priori knowledge of cutoffs. Consider a program $j$ with cutoff $c_{j}$. Suppose that students tend to prefer programs of higher quality, consistent with what is found in the literature. If applicants behave strategically, one would expect applications to program $j$ to peak among students with application scores close to $c_{j}$. If cutoffs tend to remain in the same neighborhood across years, students with application scores much higher than $c_{j}$ can expect to be admissible to more selective, higher-quality programs than $j$, which they prefer over $j$. Hence, we expect very few of these students to include $j$ on their list. As application scores drop and are closer to $c_{j}$, students’ chances of admission to the most selective programs decrease, and program $j$ becomes one of the most selective (desirable) programs among those for which they still have a high admission probability. Hence, we expect applications to $c_{j}$ to increase as application scores decrease and draw closer to $c_{j}$. As application scores decrease below $c_{j}$, students realize that their probability of admission to program $j$ is lower, and while program $j$ remains a relatively desirable (selective) alternative, we expect these expectations to drive applications down. This application pattern, expected if students behave strategically, is exactly what we observe in the top panel of Figure 2. Pooling all programs $j$ together, the top panel of Figure 2 shows the fraction of students listing program $j$ in their rank-order lists (ROLs or $P$ in terms of our notation), as a function of the distance between their priority score for program $j$ and the cutoff $c_{j}$.
@@
It may be difficult to disentangle the role of preferences from the role of expectations about admission probabilities when both may enter students’ choice of which programs to include in their ROLs (Manski (2004); Agarwal and Somaini (2018)). The pattern observed in the top panel of Figure 2 could, alternatively, be consistent with students not behaving strategically but preferring programs that are a good fit in terms of quality, that is, programs in which their skill level would be close to the average skill level. If this were the case, applications should peak among students whose skills (proxied by application score) are close to the mean skill level in the program. This is not what we observe in the bottom panel of Figure 2. Pooling all programs $j$ together, the bottom panel of Figure 2 shows the fraction of students including program $j$ in their list, as a function of the distance between their priority score for program $j$ and the mean application score among students admitted to $j$. Conditional on application score, the share of students applying to a program $j$ is not the highest for application scores close to the mean score among students admitted to $j$. It peaks well below this level, showing that students do not systematically prefer programs in which they would be the “average” student. This is again consistent with the hypothesis that students behave strategically.
@@
This figure provides evidence of strategic behavior in student applications. Pooling all programs $j$ together, the top panel of the figure shows the fraction of students listing program $j$ in their rank-order lists (ROLs or $P$ in terms of our notation), as a function of the distance between their priority score for program $j$ and the cutoff $c_{j}$. Pooling all programs $j$ together, the bottom panel shows the fraction of students including program $j$ in their list, as a function of the distance between their priority score for program $j$ and the mean application score among students admitted to $j$.